# Onion Routing Key Distribution for QKDN

### v1

@@ I Introduction
In the case of symmetric encryption algorithms such as AES and cryptographic hash functions such as SHA-2 and SHA-3, the quantum threat is also relevant, although less severe. Grover’s algorithm can reduce the time required for an exhaustive search (brute force) attack from $2^{n}$ to $2^{n/2}$, which implies that, to maintain the same level of current security, the size of the keys would have to be doubled. In other words, to achieve the current level of security provided by AES-256, AES-512 would need to be employed.
@@ I-A Response to the quantum threat
Nevertheless, the implementation of QKD solutions depends on the development of quantum hardware, which represents a significant challenge. Although feasible, the current technology is expensive and requires direct fiber optic connections without repeaters, limiting its applicability to specific scenarios. Furthermore, QKD does not replace classical cryptography in its entirety, but rather focuses on improving the security of key exchange without providing solutions for asymmetric cryptography or cryptographic signatures.
@@ III Proposal of QKDN protocol
From the information presented in the previous sections, it is evident that both Post-Quantum Cryptography (PQC) and Quantum Key Distribution (QKD) have their own advantages and disadvantages. While PQC is notable for its versatility and applicability in conventional environments, QKD provides a superior level of security. In light of this dichotomy, the scientific community has initiated the development of hybrid solutions that integrate both technologies, thereby optimizing both security and implementation feasibility. A notable proposal entails the incorporation of PQC signature algorithms, such as Dilithium, to authenticate QKD nodes in key-relay and TN models[15]. This improvement enables the verification of messages transmitted through classical channels, which facilitates the detection of Man-in-the-Middle (MitM) attacks and prevents spoofing attempts.
@@
Nonetheless, uncertainty remains with regard to the protection of the shared secret from the intermediate nodes in both QKDN models analyzed (key-relay and Trusted Node). Just as PQC signatures for message authentication, PQC algorithms, in this case public key kind such as Kyber, can also be employed to ensure the confidentiality of messages transmitted by QKD nodes over a classical channel. If PQC encryption were applied correctly, it would prevent intermediate QKD nodes from being able to obtain the shared secret between the end nodes during their transmission.
@@
In the key-relay scheme, the implementation of PQC public key encryption is relatively simple. The initiating node encrypts the shared secret $S$ with the destination node’s PQC public key, keeping the rest of the procedure unchanged. Thus, when an intermediate node decrypts the ciphertext with the QKD key shared with its previous node, it does not get $S$ in plaintext, but $S$ encrypted with the destination node’s public key that can only be decrypted with the destination node’s private key. Finally, the destination node decrypts the received ciphertext with its QKD key and, subsequently, with its PQC private key to recover $S$.
@@
On the other hand, in the TN model, the incorporation of PQC public key is also straightforward. Rather than transmitting its cryptographic material in clear form, each node employs encryption with the TN’s PQC public key. Subsequently, the TN, upon receiving the encrypted messages, employs decryption with its PQC private key, calculates the ciphertext $C$, encrypts it with the destination node’s PQC public key, and forwards it. The destination node, upon receiving $C$, employs its private key PQC to decrypt it and thereby obtain the shared secret.
@@
In terms of performance, the incorporation of PQC public key results in increased latency in the TN architecture relative to key-relay. In key-relay, a single encryption and decryption process with PQC is sufficient, whereas in TN, this procedure must be repeated for each node. If the TN corresponds to the end node, one iteration can be eliminated; however, in scenarios involving multiple intermediate nodes, this optimization is negligible. Regardless of the QKD network model selected, hybridization with PQC public key encryption provides the robustness of QKD key-based encryption along with the flexibility of post-quantum cryptography. Therefore, in the event of an intermediate node in the QKD network being compromised, the shared secret will remain protected and will not be exposed. Conversely, in the event of a vulnerability in the PQC algorithm employed, the shared secret remains inaccessible to external attackers and is exclusively discernible by the QKD nodes within the network, thus preserving the security of the original system without hybridization.
@@ III-A Proposed protocol
Figure 4: Operation diagram for the proposed protocol to ensure communication of the secret between two nodes within a QKDN while maintaining confidentiality, integrity, authentication and anonymity. In the diagram the keys $(pk_{P_{i}},sk_{P_{i}})$ represent the PQC encryption key pair.
The hybrid QKD model with PQC signature and public cryptography almost ensures compliance with the principles of the CIA triad (Confidentiality, Integrity and Authenticity), only integrity needs to be guaranteed. It is true that integrity may be implicit in the model since commercial QKD nodes, such as Cerberis XGR QKD by ID Quantique [16], already guarantee integrity between pairs of nodes in the key sharing process, but not in external messages to the QKD protocol.
@@
    Each node will have a key pair consisting of a public key and a private key to encrypt/decrypt messages base in a PQC algorithm, e.g. Kyber.
@@
In the following sections the steps to create and processing the ciphertexts, onions, of the QSNS protocol proposal, based on [14], will be described. For this scenario, it is assumed that all the above-mentioned assumptions are met and that the initiator node has previously obtained, in a secure manner, the public keys of the other nodes in the network. For a better understanding, Figure 4 shows the messages exchanged in the protocol and Table I shows the used nomenclature where $D^{\text{sym}}(k,E^{\text{sym}}(k,m))=m$, $D^{\text{asym}}_{\mathsf{sk}}(E^{\text{asym}}_{\mathsf{pk}}(m))=m$ and $\operatorname{Ver}_{pk_{i}}(t,\operatorname{Sig}_{sk_{i}}(t))=1$. Additionally, $(pk_{i},sk_{i})$ for encryption and signing have the same nomenclature but are different key pairs.
@@
$E_{\mathsf{pk}_{i}}^{\text{asym}}(\cdot)$ Public key $\mathsf{pk}_{i}$ encryption with Kyber
$D_{\mathsf{sk}_{i}}^{\text{asym}}(\cdot)$ Private key $\mathsf{sk}_{i}$ decryption with Kyber
$E^{\text{sym}}(k,\cdot)$ Symmetric key $k$ encryption with AES-512-CBC
$S^{\text{sym}}(k,\cdot)$ Symmetric key $k$ decryption with AES-512-CBC
$\mathsf{Sig}_{\mathsf{sk}_{i}}(\cdot)$ Sign with Dilithium
$\mathsf{Ver}_{\mathsf{pk}_{i}}(\cdot,\cdot)$ Verify sign with Dilithium
@@
    Compute original onions: $O_{n+1}=(P_{n+1}||E^{\text{asym}}_{\mathsf{pk}_{n+1}}(S))$ and $O_{i}=(P_{i}||E^{\text{asym}}_{\mathsf{pk}_{i}}(O_{i+1}))$ for $i=n,\dots,1$.
@@
        $\tau_{i}\longleftarrow\mathsf{Sig}_{CD}(sk_{i},(O_{i+1}||B^{\prime}_{1}||\cdots||B^{\prime}_{N-1}))$
@@
        $B_{\text{new}}\leftarrow E^{\text{asym}}_{\mathsf{pk}_{i}}(\mathsf{embed}(k_{i},\tau_{i}))$
@@
      $B_{1},\dots,B_{N})\leftarrow(B_{\text{new}},B^{\prime}_{1},\dots,B^{\prime}_{N-1}).$
@@
    Then, it encrypts $O^{\prime}_{1}=(O_{1}||t_{1}||\mathsf{rdm}_{1})$ with the QKD key $k_{P_{0},P_{1}}$ [^2] The algorithm for this symmetric encryption will be AES achieving semantic security as OTP is not feasible due to the length of the delivered message. and sends the ciphertext to $P_{1}$ [^3] It is not necessary to sign the $O^{\prime}_{i}$ ciphertext since they are encrypted with a symmetric QKD key and when these are negotiated they already guarantee the integrity of the nodes..
@@
    Decrypt the incoming ciphertext with $k_{P{i-1},P_{i}}$ getting the onion $O^{\prime}_{i}$.
@@
    Extract symmetric key and tag. $(k_{i},t_{i})\longleftarrow\mathsf{Extract}\bigl{(}D^{\text{asym}}_{\mathsf{sk}_{i}}(B_{1})\bigr{)}$
@@
    Then, encrypt $(P_{i+1},O^{\prime}_{i+1})$ with QKD key $k_{Pi,P_{i+1}}$ and send it to $P_{i+1}$.

### v2

@@ I Introduction
In the case of symmetric encryption algorithms such as AES and cryptographic hash functions such as SHA-2 and SHA-3, the quantum threat is also relevant, although less severe. Grover’s algorithm can reduce the time required for an exhaustive search (brute force) attack from $2^{n}$ to $2^{n/2}$, which implies that, to maintain the same level of current security, the size of the keys would have to be doubled. In other words, to achieve the current level of security provided by AES-128, AES-256 would need to be employed.
@@ I-A Response to the quantum threat
Nevertheless, the implementation of QKD solutions depends on the development of quantum hardware, which represents a significant challenge. Although feasible, the current technology is expensive and requires direct fiber optic connections without repeaters, limiting its applicability to specific scenarios. Furthermore, QKD does not replace classical cryptography in its entirety, but rather focuses on improving the security of key exchange without providing solutions for asymmetric cryptography, to compute shared keys via KEMs (Key Encapsulation Mechanism), or cryptographic signatures.
@@ III Proposal of QKDN protocol
From the information presented in the previous sections, it is evident that both PQC and QKD have their own advantages and disadvantages. While PQC is notable for its versatility and applicability in conventional environments, QKD provides a superior level of security. In light of this dichotomy, the scientific community has initiated the development of hybrid solutions that integrate both technologies, thereby optimizing both security and implementation feasibility. A notable proposal entails the incorporation of PQC signature algorithms, such as Dilithium, to authenticate QKD nodes in key-relay and TN models[15]. This improvement enables the verification of messages transmitted through classical channels, which facilitates the detection of Man-in-the-Middle (MitM) attacks and prevents spoofing attempts.
@@
Nonetheless, uncertainty remains with regard to the protection of the shared secret from the intermediate nodes in both QKDN models analyzed (key-relay and Trusted Node). Just as PQC signatures for message authentication, PQC algorithms, in this case public key KEM such as Kyber, can also be employed to ensure the confidentiality of messages transmitted by QKD nodes over a classical channel. If PQC encryption were applied correctly, it would prevent intermediate QKD nodes from being able to obtain the shared secret between the end nodes during their transmission.
@@
In the key-relay scheme, the implementation of PQC public key encryption is relatively simple. The initiating node encrypts the shared secret $S$ with the destination node’s symmetric key computed previously with a PQC-KEM, keeping the rest of the procedure unchanged. Thus, when an intermediate node decrypts the ciphertext with the QKD key shared with its previous node, it does not get $S$ in plaintext, but $S$ encrypted with the symmetric key shared by destination and initiator. Finally, the destination node decrypts the received ciphertext with its QKD key and, subsequently, with its PQC-KEM key to recover $S$.
@@
On the other hand, in the TN model, the incorporation of PQC-KEM is also straightforward. Rather than transmitting its cryptographic material in clear form, each node employs encryption with a symmetric PQC-KEM key shared with the TN. Subsequently, the TN, upon receiving the encrypted messages, employs decryption with the corresponding PQC-KEM key, calculates the ciphertext $C$, encrypts it with symmetric key shared with the destination node, and forwards it. The destination node, upon receiving $C$, employs the PQC-KEM key that shares with TB to decrypt it and thereby obtain the shared secret.
@@
In terms of performance, the incorporation of PQC-KEM results in increased latency in the TN architecture relative to key-relay. In key-relay, a single encryption and decryption process with PQC is sufficient, whereas in TN, this procedure must be repeated for each node. If the TN corresponds to the end node, one iteration can be eliminated; however, in scenarios involving multiple intermediate nodes, this optimization is negligible. Regardless of the QKD network model selected, hybridization with PQC-KEM encryption provides the robustness of QKD key-based encryption along with the flexibility of PQC. Therefore, in the event of an intermediate node in the QKD network being compromised, the shared secret will remain protected and will not be exposed. Conversely, in the event of a vulnerability in the PQC algorithm employed, the shared secret remains inaccessible to external attackers and is exclusively discernible by the QKD nodes within the network, thus preserving the security of the original system without hybridization.
@@ III-A Proposed protocol
Figure 4: Operation diagram for the proposed protocol to ensure communication of the secret between two nodes within a QKDN while maintaining confidentiality, integrity, authentication and anonymity. The signature keys are used in the process of creating and verifying $t_{P_{i}}$.
The hybrid QKD model with PQC signature and PQC-KEM almost ensures compliance with the principles of the CIA triad (Confidentiality, Integrity and Authenticity), only integrity needs to be guaranteed. It is true that integrity may be implicit in the model since commercial QKD nodes, such as Cerberis XGR QKD by ID Quantique [16], already guarantee integrity between pairs of nodes in the key sharing process, but not in external messages to the QKD protocol.
@@
    Each node will have a key pair consisting of a public key and a private key to encrypt/decrypt messages base in a PQC-KEM algorithm, e.g. Kyber.
@@
The following sections details the steps to create and processing the ciphertexts, onions, of the QKDN protocol proposal, based on [14]. For this scenario, it is assumed that all the above-mentioned assumptions are met, that the initiator node has previously computed a PQC-KEM key with the other nodes in the network and all the public keys were distributed among the nodes. For a better understanding, Figure 4 shows the messages exchanged in the protocol and Table I shows the used nomenclature where $D^{\text{sym}}(k,E^{\text{sym}}(k,m))=m$, and $\operatorname{Ver}\bigl{(}pk_{i},\bigl{(}t,\operatorname{Sig}(sk_{i},t)\bigr{)}\bigr{)}=1$. Additionally, $(pk_{i},sk_{i})$ for encryption and signing have the same nomenclature but are different key pairs.
@@
$E^{\text{sym}}(k,\cdot)$ Symmetric key $k$ encryption with AES-256-CBC
$D^{\text{sym}}(k,\cdot)$ Symmetric key $k$ decryption with AES-256-CBC
$\mathsf{Sig}(\mathsf{sk}_{i},\cdot)$ Sign with Dilithium
$\mathsf{Ver}\bigl{(}\mathsf{pk}_{i},(\cdot,\cdot)\bigr{)}$ Verify sign with Dilithium
@@
    Compute original onions: For the first encryption in the encapsulation $O_{n+1}=(P_{n+1}||E^{\text{sym}}(\mathsf{k}^{PQC}_{P_{0},P_{N}},S))$ and $O_{i}=\bigl{(}P_{i}||E^{\text{sym}}(\mathsf{k}^{PQC}_{P_{0},P_{i}},O_{i+1})\bigr{)}$ for $i=n,\dots,1$.
@@
        $\tau_{i}\longleftarrow\mathsf{Sig}(sk_{i},(O_{i+1}||B^{\prime}_{1}||\cdots||B^{\prime}_{N-1}))$
@@
        $B_{\text{new}}\leftarrow E^{\text{sym}}(\mathsf{k}_{i},\mathsf{embed}(k_{i},\tau_{i}))$
@@
      $(B_{1},\dots,B_{N})\leftarrow(B_{\text{new}},B^{\prime}_{1},\dots,B^{\prime}_{N-1}).$
@@
    Then, it encrypts $O^{\prime}_{1}=(O_{1}||t_{1}||\mathsf{rdm}_{1})$ with the QKD key $k^{QKD}_{P_{0},P_{1}}$ [^2] The algorithm for this symmetric encryption will be AES achieving semantic security as OTP is not feasible due to the length of the delivered message. and sends the ciphertext to $P_{1}$ [^3] It is not necessary to sign the $O^{\prime}_{i}$ ciphertext since they are encrypted with a symmetric QKD key and when these are negotiated they already guarantee the integrity between a pair of nodes..
@@
    Decrypt the incoming ciphertext with $k^{QKD}_{P_{i-1},P_{i}}$ getting the onion $O^{\prime}_{i}$.
@@
    Extract symmetric key and tag. $(k_{i},t_{i})\longleftarrow\mathsf{Extract}\bigl{(}D^{\text{sym}}(\mathsf{k}_{1,i},B_{1})\bigr{)}$
@@
    Then, encrypt $(P_{i+1},O^{\prime}_{i+1})$ with QKD key $k^{QKD}_{P_{i},P_{i+1}}$ and send it to $P_{i+1}$.